Fermi National Laboratory


Accelerator Update

Monday June 23
The day shift began with Operations monitoring store 2715, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

A Tevatron (TeV) kicker power supply (C:PAK5PS) failed at 8:20 AM. This kicker was needed to terminate the colliding beams. Experts are investigating other options.

Two Booster RF Stations (BRF7 & 17) tripped off twice between 8:30 and 9 AM.

A MiniBooNE quadrupole power supply (E:Q874) tripped off at 10:53. There was no beam to MiniBooNE at the time of the trip. Experts investigated and found a controls problem. They had the supply on by 11:51 AM.

Booster experts began tuning the Booster beam at 11 AM.

Linac experts began tuning their machine at 12:18 PM.

MiniBooNE's Q847 tripped off at 2:22 PM. An expert investigated and found a loose connector in its breaker panel.

Operations terminated store 2715 at 4:30 PM.

The Collider Detector Facility (CDF) reported having a power supply failure that required an immediate access. Operators began preparing the enclosure for an access. Both CDF and the D-zero (D0) experimenters were accessing their collision halls by 7 PM.

The Main Injector's RF region suffered a vacuum trip at 7:51 PM.

D0 completed their access at 10:51 PM. CDF came out of access at 11:21 PM.

Operations put the TeV through a dry squeeze at 11:37 PM.

Tuesday June 24
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

Booster suffered trips of two RF stations (BRF4 & 13) at 12:17 AM. The trip indication was a power amplifier LCW leak. Operators reestablished beam 15 minutes later.

Operations established store 2718 at 2:15 AM with an initial luminosity of 39.4E30. Pbar resumed stacking.

At 7:42 AM, operators began preparing the Linac, Booster, MI, and MiniBooNE tunnels for a scheduled access.

The Linac access ended at 9:30 AM. MiniBooNE was out of access by 10:09 AM.

Operators established beam to the Booster dump at 11:40 AM.

The MI tunnel was made up by 12:15 PM.

Operations established beam to MiniBooNE at 1:08 PM. Problems developed with the MiniBooNE multiwires. Experts issued a "Set" command and resolved the problem around 2 PM.

Operations terminated store 2718 at 10:08 PM.

Vacuum technicians began working on the TeV sector E1 controls at 10:55 PM. At 11:30 PM, the vacuum experts were still trouble shooting problems.

Operations put the TEV into shot setup at 11:43 PM.

Wednesday June 25
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

The TeV suffered a sector BA quench at 1:36 AM, which caused the loss of store. Pbar resumed stacking. Experts investigated.

Operations began a dry squeeze at 5:57 AM, and the TeV suffered another sector BA quench on the first TeV ramp.

Today's Plan
The plan for today is to allow TeV experts to investigate the causes for the two TeV quenches at sector BA. Shot setup is tentatively planed for this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/25/2003   email Fermilab